OpenStudy (anonymous):

The function f(x)= x ln(x^a)-ax is defined for some positive constant a. For what x-value(s) does f'(x)=0?

OpenStudy (anonymous):

can you get \[f \prime(x)\]

OpenStudy (anonymous):

ln(x^a)+xaln(x)-a?

OpenStudy (anonymous):

\[ \ln(x^a) = a\ln(x) \]

OpenStudy (anonymous):

no :( \[y=xlnx ^{a}-ax\] \[y=axlnx -ax\] \[y \prime=a*lnx +ax*\frac{ 1 }{ x }-a\]

OpenStudy (anonymous):

\[y \prime=alnx\]

OpenStudy (anonymous):

now you have to \[y \prime=0\]

OpenStudy (anonymous):

\[alnx =0\]

OpenStudy (anonymous):

Isn't it f'(1)=0

OpenStudy (anonymous):

yes ln1=0